What would be the dimension of a rectangle if the area is 12 and the perimeter is 16?
Answer by checkley77(12844)

12=LW OR L=12/W FOR THE AREA.
2L+2W=16 FOR THE PERIMETER.
2(12/W)+2W=16
24/W+2W=16
(24+2W*W)/W=16
(24+2W^2)/W=16
24+2W^2=16W
2W^2-16W+24=0
2(W^2-8W+12)=0
2(W-6)(W-2)=0
W-6=0
W=6 ans. L=12/6=2 ANS.
W-2=0
W=2 ANS. L=12/2=6 ANS.
PROOF:
2*6=12
12=12
2*2+2*6=16
4+12=16
16=16